Session: Feb 2012 (session 1)
Wavelength: 5 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Roda at Fri Feb 24 11:48:11 2012) Minor failures: FS
computer problems. Lost scan from N80 to N90

Torun (Kaz Borkowski at Fri Feb 24 11:50:40 2012) Minor failures:
Lost 7:12 to 7:25 due to control system hang-up. No other known problems.

Westerbork (Geert Kuper at Fri Feb 24 12:11:01 2012) Success:

Yebes (Daniel Sastre at Fri Feb 24 19:17:20 2012) Success:
No known problems.

Onsala (Michael Lindqvist at Sat Feb 25 08:17:36 2012) Minor failures:
Lost several hours due to wind.

Effelsberg (Uwe Bach at Mon Feb 27 08:48:54 2012) Minor failures: ANT
Tracking problems due to control PC. Not on source from 8:45 to 9:30 UT.

Jodrell_Bank (Alastair Gunn at Mon Mar 19 15:11:50 2012) Success:
No known problems.